# Generalizability Under Sensor Failure: Tokenization + Transformers Enable More Robust Latent Spaces

### 28 Feb 2024 v1

@@ 1 Introduction
In nearly all neural datasets from microelectrode arrays to scalp electroencepholography (EEG) signals, we collect recordings from many sessions and subjects, and would wish to be able to model all the data togther in common latent spaces for further analysis. Prior work has begun to study generalizability along the session and subject dimensions (Peterson et al. (2021); Talukder et al. (2022)), but few study robustness to sensor failure which is highly prevalent in neuroscience experiments Talukder et al. (2022). Sensors can fail or become noisy within a session and sensor counts can vary between recording setups, such as EEG caps being recorded with different number of channels. Instead of reducing the dataset down to the common clean channels, which is the primary technique of dealing with noisy or missing sensors, it would be beneficial to be able to have a model robust to sensor failure.
In this work, we systematically study two time series models and their generalizability across sessions, subjects, and sensor failure levels. We introduce a framework to test sensor failure robustness (Figure 1), collect a new dataset that bolsters our ability to study these generalization cases, evaluate two models’ ability to generalize, and explore the latent space of the more generalizable model. The two models we compare include EEGNet (Lawhern et al. (2018)), the most used prior work which leverages a convolutional neural network (CNN) upon which many other models (Peterson et al. (2021)) are built, and TOTEM (Tokenized Time Series Embeddings) (Talukder et al. (2024)), the most recent prior work which leverages tokenization + transformers. It was recently found that TOTEM exhibits strong generalizability capabilities within data modalities on various time series datasets (Talukder et al. (2024)), and combined with its flexible transformer architecture, we hypothesized it could do well on additional generalizability cases such as sensor failure. We find that tokenization + transformers outperform on generalization cases and have generalizable latent spaces ripe for analysis.
Figure 1: Generalizability testing framework. Cross subject with sensor failure example.
@@ 2.1 Experimental design
To study cross session, subject, and sensor availability cases, we designed and collected our own scalp electroencepholography (EEG) dataset featuring high channel and trial count within a session. We chose to use EEG due to its high variance along these generalization cases, allowing us to study generalizability in one of the most difficult data modalities. To study sensor failure of our models at a higher resolution, we collected 128 channels of EEG simultanously. To our knowledge, the publicly available EEG datasets rarely exceed 64 channel count (Stieger et al. (2021); Kaya et al. (2018); Tangermann et al. (2012)). Additionally, given EEG’s inherently noisy nature, we wanted larger trial count sessions to provide each model with the best chance of learning useful features, so we recorded 600 trials per session, where most datasets we found had fewer than a couple hundred trials per session.
@@
Figure 2: (a) Dataset with generalizability cases illustrated. (b) Experiment setup. Adapted from García-Murillo et al. (2023)
@@ 2.2 Models
EEGNet from Lawhern et al. (2018) is a popularly used convolutional neural network (CNN) for EEG decoding. It consists of temporal and spatial CNNs that learn kernels swept across these dimensions. Its design mimics a performant EEG processing pipeline, the Filter-Bank Common Spatial Pattern (FBCSP) (Ang et al. (2012)), learning temporal kernels that extract frequency features which then get weighted by learned spatial filters (Figure 3). EEGNet takes a fixed [Sensors x Time] matrix as input.
@@ 3.2 Latent Codebook Analysis
4 Conclusion
We find that tokenization + transformers are a promising approach to modeling time series neural data which have high variability in datasets and recordings. Specifically, we demonstrate that compared with one of the most performant and popular CNN models, tokenization+transformers outperform in numerous generalization cases. These models are also ripe for interpretability analysis which can uncover new findings about time series neural data. Our study further shows how important it is to consider generalizability cases when selecting models. Applications of this work could enable automatic noisy sensor detection, interpretability of neural signals in the temporal dimension from the temporal and spatial transformers, and building of foundation models for neural time series.

### 29 Feb 2024 v2

@@ 1 Introduction
Neuroscience experiments vary across numerous dimensions including sessions, subjects and senors (Gonschorek et al., 2021; Saha & Baumert, 2020; Parvizi & Kastner, 2018). Given this inherent variability, models with strong generalization properties are desirable. Model generalizability refers to a model’s zero shot capabilities, or the model’s ability to operate on test datasets unseen at training time. Prior work studies generalizability along the datasets’ session and subject dimensions (Peterson et al., 2021; Talukder et al., 2022; Krumpe et al., 2017), but few study model generalizability under sensor variability. Sensors vary primarily because of sensor failure and sensor count differences across recording sessions. Common practice is to reduce the train and test sets to the intersection of clean and available sensors then train models. However, this throws away difficult-to-obtain neural data. It is therefore valuable to build models that can train on all available data and infer under any type of sensor variability.
To probe generalizability across sessions, subjects, and sensors, we systematically study two time series models: EEGNet (Lawhern et al., 2018) and TOTEM (Talukder et al., 2024). EEGNet is a popular convolutional neural network upon which many other models are built or compared against (Peterson et al., 2021; Xu et al., 2021). EEGNet intakes a Sensor x Time array and applies convolutional kernels along the temporal and spatial dimensions, a common approach inspired by Filter-Bank Common Spatial Patterns (FBCSP) Ang et al. (2012). TOTEM first learns a sensor-agnostic set of discrete tokens via a self-supervised vector quantized variational autoencoder, then uses these pretrained tokens as the input to a transformer classifier. TOTEM exhibits strong generalization on numerous time series datasets (Talukder et al., 2024). To study these models’ generalizability across experimental conditions we create a taxonomy of generalization cases that encompass session, subject, and sensor variability. We collect a rich electroencephalography dataset that permits testing of these generalization cases. See Figure 1 for a visualization of cross subject generalizability under sensor failure. Finally we explore TOTEM’s latent discrete codebooks, as TOTEM demonstrates the best performance across experiments in our generalization taxonomy.
Figure 1: Overview. TOTEM and EEGNet train on data from subject A1, which has no failed sensors. Both TOTEM and EEGNet are then tested on subject B2 with artificially failed sensors. This is an example of cross subject generalizability under sensor failure.
@@ 2.1 Experimental design
We first define a taxonomy of a baseline and generalization cases, the cases are as follows: Baseline Case: Within Session - train and test on the same session. Generalization Cases: (1) Cross Session - train on one session, and test on a separate session from the same subject. (2) Cross Subject - train on one subject, and test on a separate subject. (3) Sensor Failure - randomly fail Xof sensors, $X\in\{0,10,20,...,100\}$. Notably the within session, cross session, and cross subject cases can be combined with the sensor failure condition, e.g. one can study cross subject generalizability under sensor failure (Figure 1). For a visualization of the baseline and generalizability cases see Figure 2(a).
We then design and collect an electroencepholography (EEG) dataset featuring high sensor count (128 electrodes) and high trial count (600 trials/session) across two human subjects. In total this leads to four sub-datasets across subjects A,B and sessions 1,2: A1, A2, B1, B2. Our 128 sensor setup introduces relatively large datasets as previous publicly available EEG datasets rarely exceed 64 sensors (Stieger et al., 2021; Kaya et al., 2018; Tangermann et al., 2012). Additionally, much prior work with more than 64 channels collects <200 trials per session (Gwon et al., 2023; Xu et al., 2020) we collect 150 trials per direction (600 trials per session). Designing our experiments to generate relatively large datasets enables our study of generalization across numerous experimental conditions.
@@
We chose to use EEG due to its high variance along these generalization cases, allowing us to study generalizability in one of the most difficult data modalities. However, challenges regarding sensor variation are highly prominent in larger signal-to-noise ratio recording modalities such as electrocorticography or multi-unit probes (Neuropixels, Utah arrays, DBS probes). In the future we aim to replicate our EEG experiments with other neural data recording modalities, see Section 4 for further discussion.
Figure 2: (a) Visualization of baseline (within session) and generalizability cases (cross subject, cross session, sensor failure) illustrated. (b) Experimental setup. We test on 2 human subjects each with 128 electrodes generating 600 trials per session. Adapted from García-Murillo et al. (2023)
@@ 2.2 Models
EEGNet from Lawhern et al. (2018) is a popularly used convolutional neural network (CNN) for EEG decoding. It consists of temporal and spatial CNNs that learn kernels swept across these dimensions. Its design mimics a performant EEG processing pipeline, FBCSP (Ang et al. (2012)), learning temporal kernels that extract frequency features which then get weighted by learned spatial filters (Figure 3). EEGNet takes a fixed [Sensors x Time] matrix as input.
@@ 3.2 Latent Codebook Analysis
4 Conclusion + Limitations + Future Work
We find that tokenization + transformers are a promising approach to modeling time series neural data which have high variability in datasets and recordings. Specifically, we demonstrate that compared with one of the most performant and popular CNN models, tokenization+transformers outperform in numerous generalization cases. These models are also ripe for interpretability analysis which can uncover new findings about time series neural data. Our study further shows how important it is to consider generalizability cases when selecting models.
Currently we only model sensor failure as zeroing out data. This masking technique is widely used machine learning, but does not cover all sensor failure cases. Creating more sensor failure modes is a meaningful direction for future work. Extending our interpretability analysis to the downstream transformer is another valuable future direction. We would also like to adapt the framework to work with more types of neural time series data such as sparsely sampled neural time series data such as LFP from neuropixels, Utah Arrays, and stereoencepholography (sEEG) recordings, and more varied layout modalities such as electrocortiography (ECoG). This work could enable automatic noisy sensor detection, interesting interpretations of temporal and spatial dimensions, and building of foundation models for neural time series.